Meeting Information

  • 1600 UT (1000 MDT)
  • AOC 317; IP: 146.88.16.8; phone: +1 575-835-7370

Agenda/Minutes

News

  • Working on outstanding issues with regard to data quality/quantity; focus on the CBE this week.
  • Kicked-off the 'triggered' testing; need to demonstrate 24 hour turn-around in analysis.
  • PASEO close-out summary (Bob, Claire)
  • D configuration extended until at last 23 August; final date by end of week.
  • Staff Travel:
    • Krauss: 23 Jul - 27 Jul
    • McMullin: 26 Jul - 30 Jul
    • van Moorsel: 26 Jul - 30 Jul
  • Visitors:
    • Bob Sault: 26 July - 27 August (?)
    • Brent Carlson: 23 August - 04 September (tentative)

Critical Commissioning review/reports

  • Overall status: Re-establish stability following last week's development; focus on CBE issues.
    • Kerry, Ray, Brent worked on bad cables, connectors; believe all connectivity from StB to XBB are good; significant number of errors in the synchronization; new FPGAs - working to re-establish the baseline operations. Okay for low frequency OSRO/RSRO (date is coming in from several projects now); high frequency RSRO still blocked by missing scans, etc.
  • Highest priority development areas; readjusted based on resolution of several outstanding OSRO blockers.:
    • 1) Commissioning tests/system/correlator stability (particularly focused missing data/data quality issues).
    • 2) RSRO/ECSO science observing. Have SBs ready for (AB1345, AS1015, AW775; test ready for AB1346 (TSPE0001.1177894)).
    • 3) 3-bit sampler testing; Full test plan needs to be completed by 20 August to maintain January timeline for production equipment. Must outline further tests soon to meet this schedule.
    • 4) OSRO science observing issues/observations,
    • 5) Demo science observing; have SBs ready for Crab, IRC+10216, M51 (pending correlator config issues; need test), M87.

System Verification/Hardware Commissioning:

Testing and characterization of the array based on the design goals.

Targets

Area Target Priority JIRA Deliver Status
High Frequency Reference pointing quantitative analysis 0 ECSV-16/ECDATA-91 2010Q3 Issues uncovered - see report below
Calibration Phase transfer 0 ECDATA-271 2010Q3 Needs better weather for real conclusions.
Calibration Switched power gain correction 0 - 2010Q3 pending SDM implementation for further work
Calibration RFI Sweep for all bands 2 - 2010Q3 Pending wide band widths
Calibration Develop set of frequency accuracy tests (X band is complete) 2 - 2010Q4 Perley:
Calibration Delay clunking 2 - 2010Q4 Rupen/Pokorny:

Reports

  • Antenna: Perley
    • Monitoring update: CXobs results
    • syscoll update (Sowinski) Any progress in understanding?
    • pn3db update (Sowinski/Dhawan)
  • Receiver: Perley
    • Ku rxs coming (2) next week. Pushed along - 21st now. S band this week (. First wide band X by the end of the month. Kas coming for 11 and 26 this week.
    • RP noted that we are identifying which Rxs have the quad hybrid; compare performance.
  • Array: Perley
    • Reference pointing post TelCal fix (Perez)
  • Correlator: Rupen
    • Succinct list of issues: (as last week): missing data; Michael reviewed all data in July and sees several different scenarios leading to missing scans; there are 4 scans that indicate mislabeled scans; both point to the CBE; still some cases of incoherence particularly following reference pointing.
      • Need decision on whether we edit existing files until fixed (van Moorsel).

Observing Band Checkout

Encompasses principally core operation at those frequencies followed by detailed calibration strategies, flagging.

Targets

Reports

  • Low Frequency (L/S/C): Momjian
    • Status: Operating; users should include one of the canonical flux calibrators in SB.
    • Upcoming:
      • Need switched power
        • Develop set of observations to check Tcal/Tsys/SEFD across bands
    • Links: ECSV-38

  • Mid Frequency (X/Ku): Dhawan
    • Status: Operating; users should include one of the canonical flux calibrators in SB.
    • Upcoming:
      • Need switched power
        • Develop set of observations to check Tcal/Tsys/SEFD across bands
    • Links

  • High Frequency (K/Ka/Q): Sjouwerman/Claussen
    • Status:
      • Operating; users should include one of the canonical flux calibrators in SB
      • Reference pointing in C or X
      • No observing at Ka < 32 GHz; No longer blocked; SSS/OPT code is consistent with limitations in hardware.
    • Upcoming:
      • Need switched power
        • Develop set of observations to check Tcal/Tsys/SEFD across bands
    • Links: ECSV-30

  • Update on Switched power

Observing Support:

Testing and development of the end-to-end capability of the array to support observations to the community (Documentation, OPT, Scheduling, Instrument models, Calibrator survey/models DB, Rapid response observing, demo science, post-processing).

Targets

Area TargetSorted ascending Priority JIRA Deliver Status
Software CASA development cycle priorities 0 -   Spreadsheet developed; time allocations needed for P1s
Software CBE averaging/decimation (time/freq) 1 - 2010Q3; vci averaging - honors other-than-1s dumps.  
Software/Calibration Recommended data path for weights in calibration 1 - 2010Q3  
Flux Calibration SBs for Flux calibrator models 1 ECSV-43 2010Q3 Needed for D config (2 freqs/band); 5 h of observations ready

Reports

  • Meta summary
    • Multi-config SBs in regular use; note 'overhead' in configuration changes; Must add 30 seconds to each configuration change - Gustaaf's group will check projects for this until relaxed, in particular, reference pointing must expand to 3 minutes..

  • Software Systems (Overlord: Butler):
    • Versioning Page: https://staff.nrao.edu/wiki/bin/view/EVLA/EVLASoftwareVersioning
    • *CBE store: New lustre file system Wednesday.
    • OPT (Lorant):
    • Scheduling (Joan):
    • Executor (Barry):
    • Corr (CBE/CM; Michael):
    • Archive (John): Status of security hole.
    • Post-processing
      • CASA (Nick):
      • AIPS (Eric):
    • ASDM
      • SDM total power/switched power table write-up (Rupen).
      • Flag table write-up (Myers/McMullin).Draft in and being reviewed; will make changes and percolate more broadly.
      • SDM document (with at least table descriptions/contents/units)Needed by the end of the month. (Rupen).
      • Priority 3: Ephemeris table resolution (Butler)
      • Bryan raised the issue of changes to the filler for more precise UVW calculations; will contact Joe on this.

  • Flux Calibrator Models (Amy/Rick): TCAL
    • Report: TCAL0001_sb1576498_1.55377.02314818287, TCAL0001.sb1575783.eb1620753.55378.450005439816. 5 hours of observations taken (low frequencies). Postpone report to next week.
    • Polarization calibrator list (Myers/Perley)
    • Links: ECSV-43

Observing Mode Support:

Testing and development within broad observing modes distinguished by their science goals (continuum, spectral line, polarization).

Targets

Area Target Priority JIRA Deliver Status
Spectral Line Bandpass stability 1 ECSV-35 2010Q3  
Spectral Line Doppler tracking 1 ECSV-32 2010Q3  
Spectral Line Spectral line setups; calibration procedures recommendations (OSRO/RSRO) 1 ECSV-31 2010Q3  
Polarization On-axis leakage calibration 1 ECSV-25 2010Q3  
Polarization R-L phase calibration 1 ECSV-25 2010Q3  
Polarization Observer Recommendations (OSRO/RSRO) 2 ECSV-25 2010Q3  
Polarization Beam squint calibration 2 ECSV-25 2010Q3  
Polarization Primary beam (off-axis) calibration 2 ECSV-25 2010Q3  
Polarization Polarization calibration service observation plan 3 ECSV-25 2010Q4  
Demo Science SBs for approved EPO projects 1 ECSV-57 2010Q3 Started for a subset

Reports

  • Spectral Line Modes (Ott/Goss): TSPE
    • Testing:
      • Following discussion last week, test SB for demonstration (Ott report).
      • Spectral Line general commissioning tests (diff spw widths, diff frequencies, bandpasses, etc): ECSV-92 * Bandpass stability: ECSV-35: More date coming soon; asked to revise source to take it out of the day time competition with development/construction activities.
    • Links: ECSV-19
  • Polarization Modes (Myers): TPOL
    • Status: Meeting minutes posted in: https://bugs.aoc.nrao.edu/browse/ECSV-24
    • Testing
      • Low frequency check outs completed; derived fractional polarization for known sources agree with predictions.
        • C band report (Marvil). ECSV-69 shows the recent results from the 3 hour C band; overall things look good and quite stable; report will be posted into JIRA.
      • Multi-band leakage/pos. angle calibration report pending.
      • Post-processing (AIPS): AH1004 data checkout (still needed)
    • Links:

  • Demo Science (Chandler/Condon): TDEM
    • Status: Begun; list requires prioritization; limited remaining time in D-config; Will try for IRC+10216 this week as possible.
    • Testing
      • Cyg A (Perley): Status
      • Crab Nebula (Frail): Data taken this weekend but bad configuration so likely not much good out of the 5 hours.
    • Links: https://staff.nrao.edu/wiki/bin/view/EVLA/DemoScience

  • Advanced Planetary: TPLA
    • Testing
      • Cassini observations obtained.; Under review by Bryan.
    • Status:
      • Ready to test this week (only non-OST for now)
  • Advanced: Solar: TSOL
    • Status: Planning/prototype testing ongoing; working with L, Ku (special Rx setups in these two bands). RP noted that Tim Bastien will be out this week; following PASEO will work on solar mode testing.
  • Advanced: Pulsar (Brisken): TPUL
  • Advanced: VLBI (Walker/Brisken): TVLB
  • Advanced: Astrometry (Mioduszewski)

  • RSRO/ECSO: TRSR
    • Status: RSRO Readiness (Cycle1) Master List
      • Need RSRO/ECSO SBs in now! Please note the following:
        • Follow the nominal guide for OSRO (e.g., ensure wrap is set for first few scans, dummy scan for each frequency change).
        • Need 2x2 minute OSRO scans at the beginning of the SB to provide additional time for the correlator to configure (priming for RSRO); it's recommended that these be at a different frequency than the science observations to enable easily distinguishing them from the science data in post processing.
        • No scans longer than 90s (or so - exact number TBD).
        • Set it up using the extended OSRO configurations
    • RSRO Science Observations (Red=recent observations requiring report; Blue=active; Black=pending):
      • AB1345 (TRSR0004): Data taken; need report before further observations.
      • AC992: AC992_sb1426692_1_000.55380.88644908565 (Chomiuk).
      • AB1346 (TRSR0005): AB1346.55382.19686146991 (Brogan/Hunter). Report available in: http://bugs.aoc.nrao.edu/browse/RSRO-3; many problems; data not viable.
      • AS1015 (TRSR0010): AS1015_sb1626830_1.55380.92802219908 (Chomiuk). Report available in: http://bugs.aoc.nrao.edu/browse/ECDATA-267; data look good; time logged from RSRO project. Comparison between OSRO/RSRO observations; update (Chomiuk).
      • AV318: Converted to OSRO and completed.
      • AM1014 (TRSR0009)
      • AT374 (TRSR0011); Need test run ASAP.
      • AK726 (TRSR0007)
      • AA330 (TRSR0003)
      • AL746 (TRSR0008): TRSR0008_Ka.55382.59489054398, TRSR0008_C.55382.62113342593; Report posted at: ECDATA-268; issues with C band RFI; missing spw; B/D not so good (didn't use set-and-remember for B/D).
      • AC982 (TRSR0002): AC982_sb1537991_1.55388.49498728009 (382 GB of love).Being reviewed.;
      • AH1006 (TRSR0006)
    • Resident Schedule: Posted on home page of twiki now!
      • 01 Jan - 16 May: Adam Deller (AC982: RSRO-4): 300A
      • 31 Jan - 31 Jul: Laura Chomiuk (AS1015/AK726: RSRO-11/RSRO-6): 200C
      • 04 Feb - 01 Jun: Jacqueline van Gorkom (AV318: RSRO-13): 200D
      • 25 Feb - 01 Jun: James Miller-Jones (AM1014: RSRO-9): 200E
      • 02 Mar - 07 Jun: Manuel Aravena (AA330: RSRO-1): 300G
      • 16 Mar - 08 Apr: Dave Green (AB1345: RSRO-2): 300G
      • 20 Apr - 30 Jun: Crystal Brogan (AB1346: RSRO-3): 200B
      • 20 Apr - 30 Jun: Todd Hunter (AB1346: RSRO-3): 200B
      • 30 Apr - 15 Jul: Joseph Lazio (AC982: RSRO-4): 266
      • 07 Jun - 16 Sep: Laura Perez (AC982: RSRO-4): 200E
      • 28 Jun - 05 Oct: Volker Heesen (AH1006: RSRO-5): 200G
      • 05 Jul - 26 Oct: Russ Taylor (AT374: RSRO-12): 301; Shannon Jaeger is also here.
      • 21 Jul - 23 Sep: Dave Green (AB1345: RSRO-2): 300G
      • 15 Sep - 15 Dec: Adam Leroy (AL746: RSRO-7): 200C
      • 01 Oct - 01 Jan: Hendrik Linz (AC982: RSRO-4): 200C
      • 01 Jan - 01 Apr: Jane Greaves (AC982: RSRO-4): 300G
      • 02 Jan - 02 Apr: Stuart Corder (AC982: RSRO-4): 300G
    • Current Residents
      • Laura Chomiuk (POC: Miller Goss; area(s): Correlator checkout; Spectral line commissioning; OSRO data validation
        • This week:
      • Laura Perez (POC: Claire Chandler); area(s): High frequency observing
        • This week:
      • Volker Heesen (POC: Michael Rupen); area(s): polarimetry
        • This week:
      • Russ Taylor (POC: Rick Perley); area(s): polarimetry
        • This week:
    • Past (and future residents)
      • Joe Lazio (POC: Claire Chandler); area(s):
      • Crystal Brogan (POC: Claire Chandler); area(s): High frequency observing; demo science
      • Todd Hunter (POC: Claire Chandler; area(S): science support software interfaces for users and high frequency observing; band/polarization checkout.
      • Manuel Aravena (POC: Chris Carilli); area(s): high frequency observing
      • Adam Deller (POC: Michael Rupen); area(s): phased array system design and implementation
      • Dave Green (POC: Sanjay Bhatnagar): area(s): algorithm tests for wideband imaging (currently away).
      • James Miller-Jones (POC: Claire Chandler; area(s): demo science use cases
      • Jacqueline van Gorkom (POC: Miller Goss; area(s): Correlator checkout; spectral line commissioning)

  • Science Support Team: Adding assignments for targeted data reviews
    • McMullin: ASDM issues, data review; testing organization
    • Myers:
    • Ott:
    • Greisen:
    • Dhawan:
    • Kogan:
    • Rupen:
    • Krauss:

Advanced Science Support:

These are the testing/development areas needed to fulfill the full capabilities of the EVLA along with additional observing modes for other branches of the community.

  • Astrometry (Mioduszewski)
    • Status: Wait for A array(?)

  • HPC
    • Status: Minutes from last meeting.
The Postprocessing Hardware group met on July 2 at 1pm.

Attending in Socorro: Elias, Golap, Greisen, Owen, Robnett, van Moorsel in Charlottesville:  Cotton,Halstead, Hatz, Hunt, Lacy


  Kumar reported that had gotten the continuum simulation to run on the cluster and has circulated his primary results to this group. James argued that all the results can be understood in terms of disk head contention and the speed of the current network, which will be improved in the EVLA cluster so that the issues Kumar sees should not be a problem. Kumar will pass scripts to James next week so that James can run a set of tests himself to determine the final parameters of the EVLA cluster to be purchased.  It looks like the final parameters can be specified in time to allow a purchase by the end of July. Hopefully James can report his recommendations at the next meeting, which will be a major milestone.

  However, an important question is whether we should make the purchase then given the delays in the WIDAR commissioning. It would be good to be able to get on with the commissioning of the cluster. That means working out the procedures for submitting and queuing jobs for the cluster and then running test cases. James suggests this might be done by using a subset of the prototyping cluster and other existing hardware. If this is possible, we might defer the final cluster purchase in order to get more for our dollars; otherwise we might want to go ahead and purchase the cluster soon so we can get started.

  We further discussed the future of this meeting. The group wants to continue the twice-a-month format. However, once the EVLA cluster parameters have been settled, possibly at the next meeting, move the focus to more observatory-wide issues of HPC. Another meeting should be set up for local issues relating to the EVLA cluster.

  Because of the PASEO meeting on July 16, we will hold our next meeting on July 23rd.
    • Links:

AOB

Misc Actions

  • [McMullin]: Confirm restriction of RSRO scan length at 90 s
  • [Owen]: Outline issue with 90s scan length restriction.
  • [McMullin/Durand]: Clean up outdated MainSaver issues.
  • [McMullin]: Review/recover metrics for Priority 0 items.
  • [Perley/McMullin]: SPW/IF/BB lexicon and mapper
  • List of known problems (can be used by OSRO commissioning team (van Moorsel)
      • Med Priority [Rupen]: X, Ku band subbands are inverted (IF 1 is labelled as IF4, IF2 as 3, IF3 as 2, IF 4 as 1). Noted that this had been tested but there were enough changes to warrant a re-do.
      • Med Priority [TBD]: Phase offsets within scans (phase jump at scan boundaries). R&L phase offset on some scans - probably L302 LO problem, needs tracking down. MR: Seen on EA07, EA25 (both RR and LL)
      • Med Priority [TBD]: Added. Phase change when frequency changes (needs to be tested).
      • Med Priority [TBD]: Anti-aliasing does not behave as advertised.
      • Low Priority [TBD]: First scan often bad.
      • Low Priority [TBD]: Phase (and amp) wobbles; long history.

-- JosephMcMullin - 19 Jul 2010
Topic revision: r2 - 2010-07-20, jmcmulli
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding NRAO Public Wiki? Send feedback